Ticket Number: HEA-NOC/20020405-02 Ticket Status: CLOSE
Ticket Type: Scheduled Ticket Source: HEA-NOC
Ticket Scope: Site Site/Line: HEAnet <==> NUIG
Ticket Owner: [log in to unmask] Problem Fixer: HEA-NOC
Ticket Opened: Problem Start: 20020410 09:00 UTC +1
Ticket Closed: 20020411 14:53:33 UTC Problem End: 20020410 09:05 UTC+1
Upgrade NUIG access (main and backup) from 20Mbs to 40Mbs
Any outages should be minimal and no longer than a minute. Upgrade is scheduled for circa 09:00
20020410 09:05 NUIG and HEAnet change pvc parameters.
This ticket will remain opened for monitoring.
20020411 Circuit is stable after monitoring period. Ticket closed.
Time to Fix:
Window of 2 hours. Total changes should take 5 minutes.
Total changes less than 30 seconds.